### v1

# Many Minds from One Model: Bayesian Transformers for Population Intelligence

Abstract
Despite their scale and success, modern transformers are almost universally trained as single-minded systems: optimization produces one deterministic set of parameters, representing a single functional hypothesis about the data. Motivated by the idea that intelligence emerge from many minds, we propose Population Bayesian Transformers (B-Trans), which transform a standard Large Language Model into a Bayesian Transformer model to supports sampling diverse yet coherent model instances from a single set of pre-trained weights.
B-Trans introduces a Bayesian-motivated posterior proxy by treating the bias-like offsets in normalization layers as stochastic variables with a Gaussian variational approximation, inducing a distribution over model behavior without the cost of training full Bayesian neural networks. Sampling from this proxy yields a set of model instances with diverse behaviors while maintaining general competence. To preserve coherence within each generation, we freeze the sampled noise at the sequence level, enforcing temporal consistency across tokens. B-Trans allows for population-level decision-making, where aggregating predictions across sampled individuals significantly enhances exploration. Experiments across zero-shot generation, Reinforcement Learning with Verifiable Rewards (RLVR), and RL without explicit labels demonstrate that B-Trans effectively leverage the wisdom of crowds, yielding superior semantic diversity while achieving better task performance compared to deterministic baselines.
Many Minds from One Model:
Bayesian Transformers for Population Intelligence
Diji Yang and Yi Zhang University of California Santa Cruz {dyang39, yiz}@ucsc.edu
1 Introduction
The predominant paradigm in Large Language Model (LLM) yang2025qwen3; liu2025deepseek deployment treats model weights as static, deterministic point estimates. While this approach effectively leverages the knowledge acquired during pre-training wang2024mmlu, it inherently limits the model’s ability to represent uncertainty. Recent studies further suggest that well-aligned models suffer from severe structural homogenization (jiang2025artificial). Under the convergence pressures of post-training alignment, such as Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF), models tend to collapse onto a narrow set of reasoning patterns. This rigidity becomes a fundamental bottleneck in complex reasoning tasks: the model struggles to explore alternative reasoning paths, locking itself into local optima yue2025does.
Current attempts to reintroduce diversity often operate at decoding time through randomness in the action space, for example via temperature-based sampling or similar entropy-promoting strategies. While increasing temperature can prevent verbatim repetition, it often acts as a surface-level perturbation. As noted by yue2025does, in reasoning tasks such sampling can lead to incoherent degeneration, yielding statistical variation without meaningful semantic progress, rather than systematically discovering novel solution paths. The core issue is that manipulating the output tokens does not alter the underlying deterministic reasoning engine. To recover diversity of thought, we argue that we must step back from a collapsed point estimate and allow the model to explore a local proxy of the posterior neighborhood around $\theta_{\text{MAP}}$, instead of relying solely on output-token randomization.
Can we achieve meaningful population diversity without abandoning the efficiency of a single model? In this work, we demonstrate the feasibility by generating many minds that can emerge from a single model. Our key insight is that normalization layers ba2016layer; zhang2019root modulate representation geometry, and that small shifts in their bias-like offsets can induce meaningfully different functional behaviors while preserving architectural stability. We therefore replace point-estimated normalization biases with a Gaussian distribution, transforming a single transformer into a compact generator of a diverse population of model instances. Each draw from this distribution corresponds to a distinct “mind” that shares the same weights and structure, yet processes information through a different normalization regime. This targeted population modeling yields diversity where it is influential, without incurring the cost of maintaining multiple full models.
To maintain logical coherence, we enforce temporal consistency by freezing the sampled offsets for the duration of each sequence. Unlike per-token noise injection (e.g., standard Dropout), which can lead to disjointed logic, our approach ensures that each generation is produced by a single, logically consistent “persona” sampled from the population. We present B-Trans as a general population formulation (Section 3) and instantiate it through normalization-layer perturbations (Section 4); this instantiation is deliberately minimal and can be extended in future work.
We validate B-Trans through three distinct experimental lenses, each probing a different facet of population intelligence. First, in zero-shot generation, we demonstrate that B-Trans effectively breaks structural homogenization, producing functionally heterogeneous outputs. Second, in Reinforcement Learning with Verifiable Rewards (RLVR) guo2025deepseek, we show that the parameter-space diversity of B-Trans enables deep exploration, traversing sparse reward landscapes more effectively than action-space baselines. Finally, in a label-free Test-Time RL (TTRL) setting, we reveal that the implicit population utilizes the wisdom of crowds. Even without ground-truth supervision, aggregating the diverse reasoning paths of the B-Trans population outperforms deterministic baselines, showing that uncertainty can act not as noise, but as a generative engine for exploration. Collectively, these results confirm that B-Trans fundamentally alters the exploratory dynamics of the model, enabling deep and population-driven reasoning at negligible computational cost. We summarize our contribution as:
2 Related Works
Bayesian Neural Networks and Ensembles.
Bayesian methods provide a principled framework for modeling weight uncertainty (blundell2015weight). Prior work has adapted Bayesian principles to neural networks for specific pragmatic objectives. For instance, kristiadi2020being demonstrated that being “a bit Bayesian” (e.g., applying Laplace approximation only to the last layer) can improve calibration and reduce overconfidence (“knowing what you don’t know”). Similarly, jing2025kalman introduced the Kalman Bayesian Transformer to support stable sequential learning and mitigate catastrophic forgetting by treating pre-trained weights as priors. While these works primarily use Bayesian tools for calibration, robustness, uncertainty quantification, or model averaging for prediction quality, B-Trans targets a generation-centric goal: we use lightweight parameter-space stochasticity to promote exploration and semantic diversity during generation. Compared to Deep Ensembles (lakshminarayanan2017simple) which scale linearly in cost, or full BNNs which are computationally intractable, B-Trans induces this population diversity solely through stochastic perturbations of normalization biases, offering a negligible alternative specifically optimized for the autoregressive nature of Transformers.
Diversity and Mode Collapse in LLMs.
Standard decoding strategies, such as Top-$k$ and Nucleus Sampling, primarily manipulate the tail of the token distribution to mitigate degeneration, but they do not address convergence in the underlying model. Recent work points out that modern instruction-tuned models exhibit severe structural homogenization, where the probability mass over-concentrates on a narrow set of reasoning patterns (jiang2025artificial). Consequently, purely stochastic decoding methods often yield variations of a single persona rather than distinct cognitive perspectives. Methods such as Self-Consistency wang2022self aggregate multiple sampled outputs, but remain constrained by the same model-level collapse. B-Trans instead diversifies behavior by sampling model instances via parameter-space perturbations, turning a single set of weights into a population that can produce more functionally diverse generations.
Exploration in Reinforcement Learning.
Efficient exploration is critical for policy improvement, especially in reasoning tasks with sparse rewards. However, the efficacy of standard action-space entropy regularization (e.g., temperature sampling) is severely compromised by the aforementioned structural homogenization.
Recent RLVR studies observe that simply increasing output entropy often forces the model into incoherent degeneration (statistical noise) rather than systematically discovering alternative solution paths yue2025does. To address this, we revisit the concept of Parameter Space Noise plappert2018parameter, originally proposed for continuous control, and adapt it to modern LLMs. By perturbing a lightweight posterior proxy through B-Trans, we encourage deep exploration over reasoning trajectories beyond surface-level token randomization.
3 Bayesian Transformer
This section formalizes a population view of a single pre-trained model via Bayesian hypothesis sampling: we draw a sequence-level parameter sample to obtain one coherent model instance, and aggregate multiple individual for population-level thinking. [^1] Throughout, “Bayesian” is used as a principled motivation for structured uncertainty and sampling, rather than a claim of exact posterior inference for billion-parameter LLMs. Section 4 then presents a minimal instantiation that perturbs normalization offsets as a computationally negligible proxy for posterior sampling.
3.1 Local Geometry and the Laplace Approximation
Let $\mathcal{D}$ be the training dataset and $\theta\in\mathbb{R}^{d}$ be the parameters of a Transformer. We view the pre-trained parameters $\theta_{\text{MAP}}$ as the Maximum A Posteriori (MAP) estimate of the posterior $p(\theta|\mathcal{D})$. While $\theta_{\text{MAP}}$ captures a high-probability configuration, it collapses the rich information contained in the geometry of the loss landscape into a single point. To recover uncertainty, we employ the Laplace Approximation, modeling the posterior neighborhood around the mode as a Gaussian:
$p(\theta|\mathcal{D})\approx\mathcal{N}(\theta;\theta_{\text{MAP}},\Sigma).$ (1)
Computing the exact covariance $\Sigma$ (the inverse Hessian) is not feasible for billion-parameter models. Furthermore, the true posterior landscape of LLMs is highly non-isotropic (li2018visualizing), and estimating its anisotropy at inference time without training data or gradients is not well-posed. However, explicitly modeling this anisotropy without access to the full training data or gradients during inference is ill-posed. For tractability, we adopt an isotropic covariance structure $\Sigma=\sigma^{2}I$. From an information-theoretic perspective, this choice represents the Maximum Entropy distribution given a constraint on the expected Euclidean distance from the $\theta_{\text{MAP}}$ estimate. Rather than risking a mis-specified diagonal approximation that might bias the model into degenerate subspaces, the isotropic assumption effectively defines a spherical trust region for exploration. Importantly, our goal is not exact Bayesian inference, but as efficient sampling from a locally valid proxy posterior, but a practical balancing theoretical motivation with engineering feasibility.
3.2 Temporal Consistency: Token-wise vs. Hypothesis Sampling
A key challenge for Bayesian sampling in autoregressive generation is handling the time dimension. Standard stochastic regularizers, such as Dropout (srivastava2014dropout), are typically applied independently at each forward pass. For a sequence $y_{1:T}$, this implies integrating over the parameters at every token step:
$p(y_{1:T}|x)\approx\prod_{t=1}^{T}\int p(y_{t}|y_{<t},x,\theta_{t})p(\theta_{t}|\mathcal{D})\,d\theta_{t}.$ (2)
We argue that Eq. (2) is misaligned with coherent multi-step reasoning: it re-samples $\theta_{t}$ at each token, which can disrupt cross-token logical consistency by changing the underlying computation mid-generation. Instead, we adopt hypothesis sampling. In the Bayesian view, a single sample $\theta\sim p(\theta|\mathcal{D})$ corresponds to one coherent hypothesis. Once sampled, this hypothesis should remain fixed for the entire sequence. The predictive distribution is:
$p(y_{1:T}|x)=\int\left(\prod_{t=1}^{T}p(y_{t}|y_{<t},x,\theta)\right)p(\theta|\mathcal{D})\,d\theta.$ (3)
We approximate Eq. (3) using Monte Carlo integration with $K$ samples. For each sample $k$, we draw a noise vector $\epsilon^{(k)}$ once and freeze it, obtaining a specific model member $\theta^{(k)}=\theta_{\text{MAP}}+\epsilon^{(k)}$. This ensures that the entire sequence $y_{1:T}$ is generated by a single consistent instance, while enabling population-level aggregation without storing multiple full sets of weights.
1class BayesianBiasWrapper($n$.Module):
2 """Augments normalization with a Bayesian latent bias."""
3 def __init__(self, base_norm, prior_mu, prior_std):
4 self.norm = base_norm
5 self.mu, self.sigma = prior_mu, prior_std # z ~ N(mu, sigma)
6 self.z_instance = None # cached per sequence
7
8 def forward(self, x): # x: [B, T, D]
9 if self.z is None:
10 B, T, D = x.shape
11 self.z = sample_normal(self.mu, self.sigma, shape=(B, 1, D), device=x.device)
12 return self.norm(x) + self.z # broadcast along T
13
14 def reset_posterior(self):
15 """Resamples the model instance (e.g., for a new prompt)."""
16 self.z = None
17
18 def apply_bayesian_transform(model, prior_mu, prior_std):
19 """Converts standard RMSNorm into Bayesian layers."""
20 for each normalization module to target: # locate layer by model.named_modules
21 replace it with BayesianNorm(module, mu, sigma)
22 return model
Figure 1: B-Trans implementation (abstracted). B-Trans operates as a plug-and-play replacement for normalization layers. For each response, it caches a latent offset per sequence and reuse it across autoregressive steps. This approximates drawing a single coherent model instance for each generation pass, enabling diverse yet temporally consistent outputs from one set of weights.
4 Implementation
Our approach is a surgical architectural modification rather than an external sampling procedure. We align with the pragmatist view that partial Bayesian structure in carefully chosen components can already induce useful functional properties (kristiadi2020being), we implement B-Trans by wrapping RMSNorm layers with a Bayesian post-normalization offset. The modified backbone remains differentiable and compatible with standard training pipelines. B-Trans operates as a plug-and-play replacement for normalization layers, ensuring that the backbone differentiable and compatible with standard training pipelines. Consequently, our method can be seamlessly integrated with existing training workflows.
4.1 Bayesian Bias Norm Wrapper
Recent LLMs typically employ RMSNorm, which is often implemented without an explicit additive bias. In B-Trans, we treat the post-normalization additive shift as a bias-like offset and inject a sequence-level latent variable $\mathbf{z}$ after normalization.
Our decision to target bias-like offsets is motivated by prior findings elhage2021mathematical; dong2021attention that biases encode “default” priors and are crucial for rank restoration, allowing us to shift reasoning dispositions without corrupting the semantic correlations stored in weights (see more discussion in Section 4.2). Concretely, for a hidden state $\mathbf{x}$, we implement:
$\mathbf{y}=\text{Norm}(\mathbf{x})\cdot\mathbf{w}+(\mathbf{b}+\mathbf{z}),\quad\mathbf{z}\sim\mathcal{N}(\mu,\sigma^{2})$ (4)
Here, $\mu$ and $\sigma$ are hyperparameters controlling the center and spread of the population diversity, and $\mathbf{b}$ can be 0 in RMSNorm. For normalization variants that include a learned scale (and optionally a bias), these parameters remain part of the deterministic backbone, while $\mathbf{z}$ provides a stochastic post-normalization offset. Sampling $\mathbf{z}$ therefore yields multiple model instances that share the same pre-trained weights but differ in their normalization regime.
4.2 Biases as Dispositional Priors.
B-Trans introduces stochasticity through additive offsets rather than through full weight perturbations. This choice is grounded in the functional role of additive shifts in Transformer mechanics. First, additive components act as input-independent defaults: while weight matrices encode correlations between features, additive shifts set baseline writes to the residual stream and influence activation thresholds before input-specific computation elhage2021mathematical. Second, additive vectors are a practical handle for semantic steering. Work on activation engineering shows that adding vectors in the residual stream can steer high-level behaviors without broadly degrading capabilities turner2023steering; zou2023representation. Third, bias-like shifts can be important for defining operating points. Recent evidence on massive activations suggests that when explicit biases are absent, models may simulate similar effects through outlier activations sun2024massive. By sampling $\mathbf{z}$, B-Trans creates a population of dispositions that explore different reasoning trajectories while sharing the same underlying knowledge in the backbone weights.
4.3 Temporal Consistency via Noise Caching
A requirement for valid variational approximation in sequential models is maintaining a consistent stochastic mask across time steps, as established in prior work on Recurrent Neural Networks gal2016dropout. Naively sampling $\mathbf{z}$ at every token step violates this principle, resulting in a temporally incoherent model that effectively shifts its functional hypothesis mid-sentence.
To adhere to this theoretical constraint within our framework, we implement a sequence-level noise caching mechanism (see Figure 1). As detailed in the BayesianBiasWrapper, the process ensures that the stochastic perturbation defines a stable reasoning persona for the duration of the generation: (1) Initialization: Upon receiving the prompt, we sample the latent variable $\mathbf{z}$ once and cache it, effectively drawing a single sample $\theta^{(k)}$ from the approximate posterior. (2) Generation: For all subsequent autoregressive steps within the same sequence, the model applies this frozen $\mathbf{z}$ to the normalization layers.
This design ensures that the stochasticity is strictly inter-sequence rather than intra-sequence. Crucially, unlike full weight-masking methods (e.g., MC Dropout), our implementation is computationally negligible, requiring only an element-wise addition of a cached tensor without the overhead of generating large binary masks. Specifically, implementing theoretically correct (frozen) MC Dropout on a 7B model would require caching around 14GB of masks per user instance, assuming pre-expand the binary masks to FP16 to align with weight precision to avoid prohibitive on-the-fly decompression overheads. In contrast, B-Trans only caches the bias noise vectors (<1MB), reducing the memory footprint by orders of magnitude while maintaining temporal consistency.
5 Experiments
Our evaluation probes B-Trans from three complementary angles. First, we test whether sampling model instances produces functional diversity in zero-shot generation rather than surface-level randomness. Second, we study exploration in Reinforcement Learning with Verifiable Rewards (RLVR), asking whether parameter-space sampling improves search under sparse rewards. Third, we examine population aggregation in a label-free Test-Time RL (TTRL) setting, where the learning signal is induced from model outputs rather than ground-truth supervision. Across all settings, our goal is to provide a minimal proof-of-concept for population-based exploration under controlled, standard protocols, rather than an exhaustive benchmark. Accordingly, we keep training and decoding setups aligned with established baselines and focus on whether B-Trans yields consistent improvements in exploration and semantic diversity.
It is important to note that while B-Trans modifies the architecture’s normalization layers, the modified backbone remains differentiable and trainable via standard backpropagation pipelines. From the efficiency and computing concern, we follow Low-Rank Adaptation (LoRA) hu2022lora training for the RLVR and TTRL tasks. In all training settings, we inject stochasticity only during rollout for exploration, while keeping the update phase deterministic by using the mean shift only, matching standard optimization stability requirements.
Figure 2: Pass@$K$ performance on MMLU-Pro across Llama-3.1 and Qwen3 families. Solid lines denote B-Trans, while dashed lines represent deterministic baselines with high-temperature sampling. While performance at $K=1$ is comparable, B-Trans exhibits superior scaling at higher $K$ values. This widening gap indicates that the induced parameter-space diversity yields valid, functional reasoning paths rather than random noise.
Figure 3: Semantic diversity on INFINITY-CHAT. B-Trans (solid) consistently yields higher embedding distances than baselines (hatched), confirming better output heterogeneity on creativity writing.
Figure 4: PCA visualization of response embeddings from Qwen3 and Llama-3.1 models. The query is “Output a hard question to humanity (super concise and short), independent of theme.” Standard sampling confines models to disjoint semantic regions (their respective “territories”). B-Trans significantly expands this scope, allowing the model to escape its default mode and even explore semantic regions typically associated with the opposing model (highlighted in red box).
5.1 Zero-Shot Diversity and Creativity
The premise of B-Trans is that different instance samples correspond to distinct functional behaviors. We evaluate whether B-Trans yields meaningful heterogeneity compared to high-temperature decoding.
Setting.
We study intrinsic population behavior on two benchmarks with different measurement goals. On MMLU-Pro wang2024mmlu (reasoning), we use Pass@K to test whether additional sampled individuals provide functional and valid reasoning paths: for each question we sample $K$ independent sequence-level personas and count the answer as correct if any sample is correct. Thus, improved scaling with larger $K$ indicates that the induced diversity is not mere noise but yields additional correct solutions. On INFINITY-CHAT jiang2025artificial (open-ended creativity), where there is no ground-truth correctness, we measure semantic diversity using average pairwise embedding cosine distance reimers2019sentence; song2020mpnet and visualize the embedding geometry with PCA following Artificial Hivemind (jiang2025artificial). The goal is to quantify and visualize semantic heterogeneity under the same generation setup as generating answer for MMLU-Pro. For both tasks, we contrast B-Trans against two distinct base models, Qwen3 yang2025qwen3 and Llama-3.1 grattafiori2024llama, using high-temperature sampling ($T=1$). This setup allows us to observe whether the models can escape their default semantic modes through B-Trans versus Temperature stochasticity.
Results.
On MMLU-Pro (Figure 2), while B-Trans matches the baseline at $K=1$, it demonstrates significantly stronger scaling behavior as the sample budget increases. This trend confirms that the diversity introduced by B-Trans is semantically meaningful: rather than generating random noise (which would yield diminishing returns), the implicit population successfully explores valid reasoning trajectories that are inaccessible to the deterministic model. Complementing this, the embedding distance analysis on INFINITY-CHAT (Figure 3) further quantifies this behavior, confirming superior output heterogeneity.
Qualitatively, Figure 4 visualizes the semantic landscape. Under standard sampling, Qwen3 and Llama-3.1 occupy disjoint, narrow regions in the semantic space, reflecting their distinct pre-trained biases. While high temperature slightly expands these clusters, it fails to alter their fundamental topology. In contrast, B-Trans generates a significantly broader distribution, forming distinct clusters that explore diverse reasoning paths.
Analysis.
The visualization in Figure 4 reveals a critical insight into the nature of the generated diversity. Standard temperature sampling is confined to local perturbations; it cannot push a model out of its inherent “personality,” leaving Qwen3 and Llama-3.1 isolated in their respective semantic territories. Conversely, B-Trans enables the models to transcend these native constraints. The visualization shows that B-Trans samples not only cover a wider area but also begin to bridge the gap between the two models, with samples from one model occasionally reaching the semantic territory typically occupied by the other. This suggests that perturbing normalization biases can shift the model’s generation distribution in embedding space beyond local temperature-based perturbations. Meanwhile, the overlap indicates that B-Trans can occasionally reach semantic regions that are atypical for the original base model under the same decoding budget, thus overlapping with the reasoning styles of different model families, rather than merely adding noise to a single persona.
5.2 Exploration in Reinforcement Learning with Verifiable Rewards (RLVR)
Model GSM8K MATH-500 Minerva Math Average
Qwen3-1.7B 86.3 47.0 21.3 51.5
      $\hookrightarrow$ B-Trans 86.8 54.4 26.1 55.7
Qwen3-8B 90.0 60.6 40.4 63.7
      $\hookrightarrow$ B-Trans 90.7 61.4 42.6 64.9
Qwen2.5-Math-7B 83.6 71.2 30.5 61.8
      $\hookrightarrow$ B-Trans 88.1 76.0 32.7 65.6
Llama3.1-8B 73.8 20.8 7.4 34.0
      $\hookrightarrow$ B-Trans 78.6 23.7 9.9 37.4
Table 1: Performance on RLVR tasks using GRPO. B-Trans consistently outperforms the baseline, particularly on smaller models and harder datasets.
In Reinforcement Learning, the critical bottleneck is often the exploration-exploitation trade-off. Standard LLM policies rely on action-space noise (sampling tokens) to explore. We hypothesize that parameter-space noise (sampling models) provides a more consistent and deeper form of exploration. By solving mathematical problems where the solution space is sparse, we test the competency of the implicit B-Trans population. If B-Trans achieves higher rewards under Group Relative Policy Optimization (GRPO) guo2025deepseek, it indicates that the diverse rollout candidates are not just different, but are hitting the correct solution regions more frequently than the baseline.
Setting.
To demonstrate the impact of B-Trans on exploration and exploitation trade-offs, we utilize the standard RLVR setting, following the SimpleRL-Zoo framework  [^2] We follow the default configuration as-is; the only change is apply BayesianBiasWrapper during the rollout, as we do not introduce any additional modifications beyond B-Trans.. To save computing, we use LoRA instead of full-parameter fine-tuning (less than 1% trainable parameters), and we train on 400 examples sampled from the same distribution of SimpleRL-Zoo dataset via difficulty-stratified sampling.
In GRPO, a group of outputs is sampled for each input, and the policy is updated based on their relative rewards. We test on four standard mathematical reasoning datasets, i.e., GSM8K lewkowycz2022solving, MATH-500 lightman2023lets, and Minerva Math lewkowycz2022solving, where the ground truth is verifiable.
Results.
We compare GRPO training with standard decoding against GRPO training where B-Trans instance sampling is enabled during rollout, under the same protocol. As shown in Table 1, B-Trans achieves consistent gains across all four benchmarks.
GRPO relies on informative relative comparisons within each sampled group. When the rollout group is homogeneous, relative advantages become less informative and updates can stagnate. By producing group members from distinct parameter-space instances, B-Trans increases the chance that at least one candidate reaches a higher-reward region, strengthening the learning signal available to GRPO under the same group size.
5.3 Test-Time RL without Labels
This experiment represents the most challenging scenario: learning without ground-truth labels. We employ Test-Time RL (TTRL), a setting where the model must self-improve by generating candidate solutions and filtering them via consistency checks. Specifically, we follow the established TTRL protocols zuo2025ttrl; yu2025restrain to utilize majority voting as the reward proxy. It is worth noting that the majority voting inherently rewards consensus, seemingly contrasting with B-Trans’s philosophy of encouraging diverse exploration (“The truth is often with the few”). However, our contribution focuses on the quality of the exploration rather than the selection mechanism. We argue that a superior learning signal should emerge from diverse independent reasoning paths rather than simple outcome consensus from mode collapse. Under the GRPO framework, standard models often generate homogeneous samples, providing redundant gradient information even when the reward is positive. Conversely, B-Trans injects parameter-space diversity, ensuring that the high-reward trajectories (those aligning with the majority) exhibit varied reasoning paths. This semantic diversity maximizes the information gain per update step, allowing the model to explore diverse reasoning patterns rather than overfitting to a single, narrow solution path.
Setting.
We investigate the efficacy of B-Trans in scenarios characterized by sparse signals and challenging search spaces using the AIME24 dataset. Unlike standard RLHF, TTRL adapts the model using only unlabeled test prompts. To synthesize a reward signal without ground truth, we utilize an outcome-based Majority Voting mechanism. For each input, we sample $N$ candidate responses and extract the final numerical answer. A consensus label is determined solely by the most frequent final answer. Trajectories leading to this consensus answer are assigned a positive reward ($r=1$), while others receive zero. Crucially, this reward ignores the semantic diversity or logical validity of the intermediate reasoning steps (Rationale), making it a coarse signal that relies heavily on the diversity of the generated candidates to be effective. All parameter updates are applied via LoRA for computational efficiency.
Results.
Figure 5 illustrates the reward curves evaluated every 5 steps. We compare the standard Qwen3-1.7B model using high-temperature sampling (temp=1.0) against B-Trans. The B-Trans exhibits high variance and oscillation but achieves significantly higher peak accuracy compared to the baseline.
Figure 5: Accuracy curve during training. Reported at every five training steps.
Analysis.
Two key insights emerge from this experiment. First, the hyperparameter $\sigma$ serves as an effective control for exploration intensity. Higher values allow the model to escape local optima through aggressive exploration, akin to brainstorming wild ideas, which is necessary for the challenging AIME24 problems. Second, the success of B-Trans in the TTRL setting suggests an emergent self-correction capability. This parallels human cognitive processes, where internal deliberation and re-evaluation can refine understanding even without external feedback. Even with a coarse consensus-based reward, B-Trans effectively filters out homogeneous cognition common to single-model instances by providing a broader set of rationales for the vote.
6 Ablation Study and Discussion
In this section, we dissect the architectural choices of B-Trans to validate our core design principles. We systematically investigate the necessity of temporal consistency in noise injection, justify the decision to perturb biases over weights, and verify the robustness of our approach across different post-training frameworks.
6.1 Temporal Consistency: Sequence-wise vs. Token-wise Noise
To validate the necessity of hypothesis sampling (Section 3.2), we compare B-Trans against a Token-wise Noise baseline on MATH-500 using Qwen3-1.8B. While B-Trans samples a single latent $z$ per sequence, the baseline resamples $z_{t}\sim\mathcal{N}(0,\sigma^{2}I)$ at every forward pass. Crucially, we fix an identical noise magnitude ($\sigma=0.02$) for both settings to isolate the impact of temporal inconsistency.
We use the Step-wise Consistency Score (SCS) to quantify the logical stability of the reasoning chain. We segment each Chain-of-Thought $Y$ into discrete steps $S=\{s_{1},\dots,s_{m}\}$ based on newline delimiters. The SCS is defined as the average cosine similarity between embeddings of consecutive steps (Encoded via all-MiniLM-L6-v2):
$\text{SCS}(Y)=\frac{1}{m-1}\sum_{i=1}^{m-1}\cos(\text{Enc}(s_{i}),\text{Enc}(s_{i+1}))$ (5)
Results & Analysis. The Token-wise baseline suffers catastrophic degradation. Despite identical noise levels, its SCS drops to 0.42 (vs. 0.58 for B-Trans), correlating with a decrease in accuracy (40.2% vs. 46.4%). This empirical failure corroborates the theoretical insight established by gal2016dropout at RNNs/LSTMs age: valid variational inference requires the stochastic mask to remain invariant across time steps to represent a single coherent hypothesis. B-Trans satisfies this “Temporal Consistency” requisite through lightweight bias perturbation, effectively capturing the benefits of weight uncertainty without the memory overhead of full ensemble masks.
6.2 Post-training Infrastructure Agnosticism
To validate that B-Trans serves as a "plug-and-play" module—agnostic not just to the model architecture but also to the training infrastructure, we repeat our TRL-based vonwerra2022trl RLVR experiments (Section 5.2) using another widely adopted RL engine, namely VeRL sheng2024hybridflow (Volcano Engine Reinforcement Learning). We observed no statistically significant difference in the final convergence performance or learning dynamics between the two implementations. While the VeRL implementation exhibited faster wall-clock training speeds in our environment, we attribute this to differences in rollout acceleration libraries rather than the B-Trans mechanism itself. We expect B-Trans to be agnostic to the underlying base LLMs and training infrastructure, supporting its potential as a surgical plug-and-play approach for diverse use cases.
7 Future Work
Our current implementation models the prior distribution as a simple isotropic Gaussian with a scalar hyperparameter $\sigma$. While this serves as an effective proof-of-concept, the assumption of uniform uncertainty across all layers and dimensions is a strong simplification.
A natural progression for this line of research is to develop a learnable or hierarchical prior. In this initial exploration, we adhered to a fixed prior to mitigate the risks associated with post-training optimization. Learning variance parameters on limited downstream datasets can lead to overfitting, where the induced population loses its general competence. We anticipate that future efforts, particularly those with sufficient computational resources to train on large-scale general corpora, could successfully learn these uncertainty parameters. Such a data-driven prior would likely act as a more robust “population,” capturing layer-specific sensitivities that a scalar $\sigma$ cannot.
Furthermore, the scalar $\sigma$ presents an opportunity for adaptive inference. The hyperparameter could be meta-learned or dynamically adjusted based on input complexity, e.g., allowing the model to be confident (low variance) on rote tasks and imaginative (high variance) on open-ended problems. Ultimately, we envision a paradigm shift where LLMs are no longer delivered as static artifacts, but as probabilistic distributions capable of adapting their internal state to the nature of the query.
8 Conclusion
We revisit the prevailing view of Large Language Models as static point estimates and argue that lightweight parameter-space uncertainty can be practically useful for exploration and diversity. We introduce B-Trans, a population-based framework that samples diverse yet coherent model instances from a single Transformer backbone using a variationally motivated Gaussian posterior proxy. By applying temporally consistent perturbations to normalization biases, B-Trans simulates the wisdom of crowds at negligible cost, yielding significant gains in zero-shot diversity and sparse-reward reinforcement learning. Our results suggest that representing uncertainty as a sampleable local proxy, even in a simple form, can enable population-driven reasoning within a single pre-trained model.
Limitations
While B-Trans demonstrates a promising direction for inducing implicit ensembles, our current study is intentionally scoped as a minimal, representative proof-of-concept centered on architectural simplicity and controlled comparisons. We do not perform extensive hyperparameter sweeps, broad architecture coverage, or large scale benchmarking beyond validating the key design principles. Our intervention is restricted to the bias terms of normalization layers, which is a simplified proxy rather than a full Bayesian neural network. While this surgical approach is computationally negligible, it is a simplified approximation of a full Bayesian Neural Network. The weights in the attention and feed-forward layers remain deterministic, potentially limiting the expressivity of the induced posterior.

### v2

@@
-Many Minds from One Model: Bayesian Transformers for Population Intelligence
+Many Minds from One Model: Bayesian-Inspired Transformers for Population Diversity
 Abstract
-Despite their scale and success, modern transformers are almost universally trained as single-minded systems: optimization produces one deterministic set of parameters, representing a single functional hypothesis about the data.
-Motivated by the idea that intelligence emerge from many minds, we propose Population Bayesian Transformers (B-Trans), which transform a standard Large Language Model into a Bayesian Transformer model to supports sampling diverse yet coherent model instances from a single set of pre-trained weights.
-B-Trans introduces a Bayesian-motivated posterior proxy by treating the bias-like offsets in normalization layers as stochastic variables with a Gaussian variational approximation, inducing a distribution over model behavior without the cost of training full Bayesian neural networks.
-Sampling from this proxy yields a set of model instances with diverse behaviors while maintaining general competence.
-To preserve coherence within each generation, we freeze the sampled noise at the sequence level, enforcing temporal consistency across tokens.
-B-Trans allows for population-level decision-making, where aggregating predictions across sampled individuals significantly enhances exploration.
-Experiments across zero-shot generation, Reinforcement Learning with Verifiable Rewards (RLVR), and RL without explicit labels demonstrate that B-Trans effectively leverage the wisdom of crowds, yielding superior semantic diversity while achieving better task performance compared to deterministic baselines.
+Despite their scale and success, modern transformers are usually trained as single-minded systems: optimization produces a deterministic set of parameters, representing a single functional hypothesis about the data.
+Motivated by the analogy to human populations, in which population-level intelligence emerges from diverse individual behaviors, we propose Population Bayesian Transformers (B-Trans), which enable sampling diverse yet coherent transformer large language model instances (hereafter referred to as a ‘mind’) from a single pre-trained LLM.
+B-Trans introduces a Bayesian-inspired posterior proxy by injecting stochasticity directly into normalization layers, avoiding the prohibitive cost of training full Bayesian neural networks.
+Sampling from this proxy yields a population of minds with diverse behaviors while maintaining general competence.
+During the generation of each response, we sample a single realization from the random distribution and hold it fixed, ensuring temporal consistency and reasoning coherence.
+Experiments on zero-shot generation and Reinforcement Learning with Verifiable Rewards (RLVR) demonstrate that B-Trans effectively leverages the stochastic model diversity, yielding superior response diversity while achieving better task performance compared to deterministic baselines.
 Many Minds from One Model:
-Bayesian Transformers for Population Intelligence
+Bayesian-Inspired Transformers for Population Diversity
 Diji Yang and Yi Zhang University of California Santa Cruz {dyang39, yiz}@ucsc.edu
 1 Introduction
-The predominant paradigm in Large Language Model (LLM) yang2025qwen3; liu2025deepseek deployment treats model weights as static, deterministic point estimates.
-While this approach effectively leverages the knowledge acquired during pre-training wang2024mmlu, it inherently limits the model’s ability to represent uncertainty.
-Recent studies further suggest that well-aligned models suffer from severe structural homogenization (jiang2025artificial).
-Under the convergence pressures of post-training alignment, such as Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF), models tend to collapse onto a narrow set of reasoning patterns.
-This rigidity becomes a fundamental bottleneck in complex reasoning tasks: the model struggles to explore alternative reasoning paths, locking itself into local optima yue2025does.
-Current attempts to reintroduce diversity often operate at decoding time through randomness in the action space, for example via temperature-based sampling or similar entropy-promoting strategies.
+The predominant paradigm in Large Language Model (LLM) Yang et al.
+(2025); Liu et al.
+(2025) deployment treats model weights as static, deterministic point estimates.
+While this approach effectively leverages the knowledge acquired during pre-training Wang et al.
+(2024), it inherently limits the model’s ability to generate diverse content, i.e., mode collapse Kirk et al.
+(2023).
+Recent studies further demonstrate that well-tuned models suffer from severe structural homogenization (Jiang et al., 2025).
+Under the convergence pressures of post-training alignment, such as Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF), models tend to collapse (converge) onto a narrow set of reasoning patterns.
+As a result, the model struggles to explore alternative reasoning paths, locking itself into local optima Yue et al.
+(2025).
+This rigidity, characterized by strong exploitation but limited exploration, can be viewed as a fundamental bottleneck in complex problem solving and learning for Artificial Intelligence Sutton et al.
+(1998).
+From the perspective of human intelligence, a population of decentralized minds can explore more insightful ideas than any single individual, a phenomenon known as the wisdom of crowds Surowiecki (2005).
+However, replicating this with a group of models is non-trivial, limited not only by high computational costs but also by the structural homogenization across model instances Jiang et al.
+(2025).
+Current attempts to reintroduce diversity at inference time operate mostly through injecting randomness into the action distribution, for example, through temperature-based sampling and related entropy-increasing strategies Wang et al.
+(2022).
 While increasing temperature can prevent verbatim repetition, it often acts as a surface-level perturbation.
-As noted by yue2025does, in reasoning tasks such sampling can lead to incoherent degeneration, yielding statistical variation without meaningful semantic progress, rather than systematically discovering novel solution paths.
+As noted by Yue et al.
+(2025), in reasoning tasks, sampling can lead to incoherent degeneration, yielding statistical variation without meaningful semantic progress, rather than systematically discovering novel solution paths.
 The core issue is that manipulating the output tokens does not alter the underlying deterministic reasoning engine.
-To recover diversity of thought, we argue that we must step back from a collapsed point estimate and allow the model to explore a local proxy of the posterior neighborhood around $\theta_{\text{MAP}}$, instead of relying solely on output-token randomization.
 Can we achieve meaningful population diversity without abandoning the efficiency of a single model?
-In this work, we demonstrate the feasibility by generating many minds that can emerge from a single model.
-Our key insight is that normalization layers ba2016layer; zhang2019root modulate representation geometry, and that small shifts in their bias-like offsets can induce meaningfully different functional behaviors while preserving architectural stability.
-We therefore replace point-estimated normalization biases with a Gaussian distribution, transforming a single transformer into a compact generator of a diverse population of model instances.
-Each draw from this distribution corresponds to a distinct “mind” that shares the same weights and structure, yet processes information through a different normalization regime.
+In this work, we demonstrate the feasibility by generating many model instances by modifying a single pre-trained model.
+Our key insight is that normalization layers Ba et al.
+(2016); Zhang and Sennrich (2019) modulate representation geometry, and that small shifts, e.g., bias-like offsets, can induce diverse behaviors while preserving semantic meaningfulness.
+We therefore replace point-estimated normalization biases with Gaussian-distributed stochastic variables, transforming a single transformer to act as a population of model instances.
+Each LLM drawn from this distribution corresponds to a distinct mind [^1] In this paper, we use the term mind interchangeably with sampled model instance.
+We adopt this terminology to emphasize that each instance acts as a distinct individual from a population.
+that shares the same weights and structure, yet processes information through a different normalization parameter (i.e.
+due to different sampled offsets).
 This targeted population modeling yields diversity where it is influential, without incurring the cost of maintaining multiple full models.
-To maintain logical coherence, we enforce temporal consistency by freezing the sampled offsets for the duration of each sequence.
-Unlike per-token noise injection (e.g., standard Dropout), which can lead to disjointed logic, our approach ensures that each generation is produced by a single, logically consistent “persona” sampled from the population.
-We present B-Trans as a general population formulation (Section 3) and instantiate it through normalization-layer perturbations (Section 4); this instantiation is deliberately minimal and can be extended in future work.
-We validate B-Trans through three distinct experimental lenses, each probing a different facet of population intelligence.
-First, in zero-shot generation, we demonstrate that B-Trans effectively breaks structural homogenization, producing functionally heterogeneous outputs.
-Second, in Reinforcement Learning with Verifiable Rewards (RLVR) guo2025deepseek, we show that the parameter-space diversity of B-Trans enables deep exploration, traversing sparse reward landscapes more effectively than action-space baselines.
-Finally, in a label-free Test-Time RL (TTRL) setting, we reveal that the implicit population utilizes the wisdom of crowds.
-Even without ground-truth supervision, aggregating the diverse reasoning paths of the B-Trans population outperforms deterministic baselines, showing that uncertainty can act not as noise, but as a generative engine for exploration.
-Collectively, these results confirm that B-Trans fundamentally alters the exploratory dynamics of the model, enabling deep and population-driven reasoning at negligible computational cost.
-We summarize our contribution as:
+We present B-Trans as a general method for sampling populations of transformer-based large language models, where each sample represents an individual mind (Section 3).
+In this paper, we develop B-Trans through normalization-layer modification (Section 4), which is deliberately minimal and can be extended to other kinds of modifications in future work.
+To maintain logical coherence, we maintain temporal consistency by using the same sampled LLM when generating a response sequence.
+Unlike per-token noise injection (e.g., standard Dropout), which can lead to disjointed logic, our approach ensures that each response is produced by a single, logically consistent model instance sampled from the population, without changing the mind in the middle of a response.
+B-Trans is validated through two distinct experimental lenses, each demonstrating different benefits of population diversity.
+In zero-shot generation, we demonstrate that B-Trans effectively breaks structural homogenization, producing semantically meaningful, heterogeneous outputs.
+In Reinforcement Learning with Verifiable Rewards (RLVR) Guo et al.
+(2025), we show that the population-based B-Trans enables deep exploration, traversing sparse reward landscapes more effectively than deterministic model baselines.
+Collectively, these results confirm that B-Trans alters the exploratory dynamics of the model, enabling deep and population-driven reasoning at negligible computational cost.
 2 Related Works
 Bayesian Neural Networks and Ensembles.
-Bayesian methods provide a principled framework for modeling weight uncertainty (blundell2015weight).
+Bayesian methods provide a principled framework for modeling weight uncertainty (Blundell et al., 2015).
 Prior work has adapted Bayesian principles to neural networks for specific pragmatic objectives.
-For instance, kristiadi2020being demonstrated that being “a bit Bayesian” (e.g., applying Laplace approximation only to the last layer) can improve calibration and reduce overconfidence (“knowing what you don’t know”).
-Similarly, jing2025kalman introduced the Kalman Bayesian Transformer to support stable sequential learning and mitigate catastrophic forgetting by treating pre-trained weights as priors.
+For instance, Kristiadi et al.
+(2020) demonstrated that being “a bit Bayesian” (e.g., applying Laplace approximation only to the last layer) can improve calibration and reduce overconfidence (“knowing what you don’t know”).
+Similarly, Jing et al.
+(2025) introduced the Kalman Bayesian Transformer to support stable sequential learning and mitigate catastrophic forgetting by treating pre-trained weights as priors.
 While these works primarily use Bayesian tools for calibration, robustness, uncertainty quantification, or model averaging for prediction quality, B-Trans targets a generation-centric goal: we use lightweight parameter-space stochasticity to promote exploration and semantic diversity during generation.
-Compared to Deep Ensembles (lakshminarayanan2017simple) which scale linearly in cost, or full BNNs which are computationally intractable, B-Trans induces this population diversity solely through stochastic perturbations of normalization biases, offering a negligible alternative specifically optimized for the autoregressive nature of Transformers.
+Compared to Deep Ensembles (Lakshminarayanan et al., 2017), which scale linearly in cost, or full BNNs, which are computationally intractable, B-Trans induces this population diversity solely through additive modification of normalization computing, offering a negligible alternative specifically optimized for modern LLMs.
 Diversity and Mode Collapse in LLMs.
-Standard decoding strategies, such as Top-$k$ and Nucleus Sampling, primarily manipulate the tail of the token distribution to mitigate degeneration, but they do not address convergence in the underlying model.
-Recent work points out that modern instruction-tuned models exhibit severe structural homogenization, where the probability mass over-concentrates on a narrow set of reasoning patterns (jiang2025artificial).
-Consequently, purely stochastic decoding methods often yield variations of a single persona rather than distinct cognitive perspectives.
-Methods such as Self-Consistency wang2022self aggregate multiple sampled outputs, but remain constrained by the same model-level collapse.
-B-Trans instead diversifies behavior by sampling model instances via parameter-space perturbations, turning a single set of weights into a population that can produce more functionally diverse generations.
+Standard decoding strategies, such as Temperature Ficler and Goldberg (2017), Top-$k$ Fan et al.
+(2018), and Nucleus Sampling Holtzman et al.
+(2019), primarily manipulate the token distribution to include diversity in model output.
+For example, methods that aggregate responses sampled via these strategies Wang et al.
+(2022) can provide slightly different outputs but still remain constrained by intra-mode collapse Kirk et al.
+(2023), i.e., a single model generates outputs with low semantic diversity due to the point-estimation nature of the underlying model.
+Another line of research improves the answer diversity in some tasks by ensembling responses from multiple LLMs (population) at a considerable computational cost Jiang et al.
+(2023).
+However, recent work points out that modern instruction-tuned models exhibit severe inter-model homogeneity Jiang et al.
+(2025), where different LLMs independently converge on similar ideas with minor variations in phrasing.
+B-Trans instead diversifies behavior by sampling model instances via parameter-space modification, turning a single set of weights into a population that can produce diverse generations.
 Exploration in Reinforcement Learning.
-Efficient exploration is critical for policy improvement, especially in reasoning tasks with sparse rewards.
-However, the efficacy of standard action-space entropy regularization (e.g., temperature sampling) is severely compromised by the aforementioned structural homogenization.
-Recent RLVR studies observe that simply increasing output entropy often forces the model into incoherent degeneration (statistical noise) rather than systematically discovering alternative solution paths yue2025does.
-To address this, we revisit the concept of Parameter Space Noise plappert2018parameter, originally proposed for continuous control, and adapt it to modern LLMs.
-By perturbing a lightweight posterior proxy through B-Trans, we encourage deep exploration over reasoning trajectories beyond surface-level token randomization.
-3 Bayesian Transformer
-This section formalizes a population view of a single pre-trained model via Bayesian hypothesis sampling: we draw a sequence-level parameter sample to obtain one coherent model instance, and aggregate multiple individual for population-level thinking.
+Balancing the tradeoff between exploration and exploitation during learning is a long-standing challenge for learning Sutton et al.
+(1998).
+Although huge efforts have been made by action-space entropy regularization during training Haarnoja et al.
+(2018); Schulman et al.
+(2017), in practice, Yue et al.
+(2025) observe that existing RL methods often force the model into incoherent degeneration (statistical noise) rather than systematically discovering alternative solution paths.
+Meanwhile, as aforementioned, empirical analysis also reveals the structural homogenization of well-aligned models, where the probability mass over-concentrates on a narrow set of reasoning patterns Jiang et al.
+(2025).
+Alternatively, instead of addressing this intractable optimization problem, we attempt to improve output diversity during the decoding stage.
+We revisit the concept of Parameter Space Noise Plappert et al.
+(2018), originally proposed for continuous control, and adapt it to modern LLMs.
+B-Trans keeps the pretrained model weight and applies modifications to a lightweight posterior proxy during model inference.
+The goal is to achieve better exploration via diverse reasoning trajectories.
+3 Bayesian-Inspired Population Model
+This section formalizes a population view of a single pre-trained model via Bayesian hypothesis sampling: we draw a sequence-level parameter sample to obtain one coherent model instance, and aggregate multiple individual models for population-level thinking.
 [^2] Throughout, “Bayesian” is used as a principled motivation for structured uncertainty and sampling, rather than a claim of exact posterior inference for billion-parameter LLMs.
@@
 3.1 Local Geometry and the Laplace Approximation
-Let $\mathcal{D}$ be the training dataset and $\theta\in\mathbb{R}^{d}$ be the parameters of a Transformer.
+Let $\mathcal{D}$ be the training dataset, and $\theta\in\mathbb{R}^{d}$ be the parameters of a Transformer.
 We view the pre-trained parameters $\theta_{\text{MAP}}$ as the Maximum A Posteriori (MAP) estimate of the posterior $p(\theta|\mathcal{D})$.
 While $\theta_{\text{MAP}}$ captures a high-probability configuration, it collapses the rich information contained in the geometry of the loss landscape into a single point.
-To recover uncertainty, we employ the Laplace Approximation, modeling the posterior neighborhood around the mode as a Gaussian:
+To recover uncertainty, we employ the Laplace Approximation, modeling the posterior proxy around the mode as a Gaussian:
 $p(\theta|\mathcal{D})\approx\mathcal{N}(\theta;\theta_{\text{MAP}},\Sigma).$ (1)
 Computing the exact covariance $\Sigma$ (the inverse Hessian) is not feasible for billion-parameter models.
-Furthermore, the true posterior landscape of LLMs is highly non-isotropic (li2018visualizing), and estimating its anisotropy at inference time without training data or gradients is not well-posed.
-However, explicitly modeling this anisotropy without access to the full training data or gradients during inference is ill-posed.
-For tractability, we adopt an isotropic covariance structure $\Sigma=\sigma^{2}I$.
+Furthermore, the true posterior landscape of LLMs is highly non-isotropic (Li et al., 2018), and estimating its anisotropy at inference time without access to training data or gradients is not well-posed.
+In principle, constructing the exact posterior geometry presents two pathways: (1) performing full Bayesian training to learn the parameter distribution from vast data associated with different LLMs (text-author pairs), or (2) computing the inverse Hessian matrix.
+However, both approaches are computationally prohibitive and data-intensive for modern LLMs.
+We therefore opt for a tractable approximation by empirically defining an isotropic covariance structure $\Sigma=\sigma^{2}I$.
 From an information-theoretic perspective, this choice represents the Maximum Entropy distribution given a constraint on the expected Euclidean distance from the $\theta_{\text{MAP}}$ estimate.
 Rather than risking a mis-specified diagonal approximation that might bias the model into degenerate subspaces, the isotropic assumption effectively defines a spherical trust region for exploration.
-Importantly, our goal is not exact Bayesian inference, but as efficient sampling from a locally valid proxy posterior, but a practical balancing theoretical motivation with engineering feasibility.
+Importantly, our goal is not exact Bayesian inference, but efficient sampling from a locally valid posterior proxy, and a practical balancing of theoretical motivation with engineering feasibility.
+While this variance is treated as a structural hyperparameter rather than a learned parameter, we show that this simplification is empirically sufficient to induce meaningful population diversity without the intractable costs of the full Bayesian alternative
 3.2 Temporal Consistency: Token-wise vs. Hypothesis Sampling
 A key challenge for Bayesian sampling in autoregressive generation is handling the time dimension.
-Standard stochastic regularizers, such as Dropout (srivastava2014dropout), are typically applied independently at each forward pass.
+Standard stochastic regularizers, such as Dropout (Srivastava et al., 2014), are typically applied independently at each forward pass.
 For a sequence $y_{1:T}$, this implies integrating over the parameters at every token step:
@@
 Once sampled, this hypothesis should remain fixed for the entire sequence.
+This also fits the real-world scenario, as each sequence is normally written by one person (which corresponds to one LLM in our sample).
 The predictive distribution is:
@@
 For each sample $k$, we draw a noise vector $\epsilon^{(k)}$ once and freeze it, obtaining a specific model member $\theta^{(k)}=\theta_{\text{MAP}}+\epsilon^{(k)}$.
-This ensures that the entire sequence $y_{1:T}$ is generated by a single consistent instance, while enabling population-level aggregation without storing multiple full sets of weights.
+This ensures that the entire sequence $y_{1:T}$ is generated by a single consistent LLM instance, while enabling population-level aggregation without storing multiple full sets of weights.
 1class BayesianBiasWrapper($n$.Module):
@@
 5 self.mu, self.sigma = prior_mu, prior_std # z ~ N(mu, sigma)
-6 self.z_instance = None # cached per sequence
+6 self.z = None # cached per sequence
 7
@@
 10 B, T, D = x.shape
-11 self.z = sample_normal(self.mu, self.sigma, shape=(B, 1, D), device=x.device)
+11 self.z = sample_normal(self.mu, self.sigma, shape=(B, 1, D))
 12 return self.norm(x) + self.z # broadcast along T
@@
 14 def reset_posterior(self):
-15 """Resamples the model instance (e.g., for a new prompt)."""
+15 """Resamples the model instance (e.g., for a new mind)."""
 16 self.z = None
@@
 19 """Converts standard RMSNorm into Bayesian layers."""
-20 for each normalization module to target: # locate layer by model.named_modules
-21 replace it with BayesianNorm(module, mu, sigma)
+20 for each normalization layer: # locate layer by model.named_modules
+21 replace it with BayesianBiasWrapper(module, mu, sigma)
 22 return model
@@
 B-Trans operates as a plug-and-play replacement for normalization layers.
-For each response, it caches a latent offset per sequence and reuse it across autoregressive steps.
+For each response, it caches a latent offset per sequence and reuses it across autoregressive steps.
 This approximates drawing a single coherent model instance for each generation pass, enabling diverse yet temporally consistent outputs from one set of weights.
 4 Implementation
-Our approach is a surgical architectural modification rather than an external sampling procedure.
-We align with the pragmatist view that partial Bayesian structure in carefully chosen components can already induce useful functional properties (kristiadi2020being), we implement B-Trans by wrapping RMSNorm layers with a Bayesian post-normalization offset.
-The modified backbone remains differentiable and compatible with standard training pipelines.
-B-Trans operates as a plug-and-play replacement for normalization layers, ensuring that the backbone differentiable and compatible with standard training pipelines.
-Consequently, our method can be seamlessly integrated with existing training workflows.
+Our approach is a surgical architectural modification during the model inference phase.
+Aligning with the pragmatist view that partial Bayesian structure in carefully chosen components can already induce useful responses (Kristiadi et al., 2020), we implement B-Trans by wrapping RMSNorm layers with a sampled offset.
+B-Trans operates as a plug-and-play replacement for normalization layers, ensuring that the modified LLM is still compatible with existing inference engines such as vLLM Kwon et al.
+(2023).
+Furthermore, B-Trans remains differentiable, which makes backpropagation possible for all kinds of learning in future works.
 4.1 Bayesian Bias Norm Wrapper
@@
 In B-Trans, we treat the post-normalization additive shift as a bias-like offset and inject a sequence-level latent variable $\mathbf{z}$ after normalization.
-Our decision to target bias-like offsets is motivated by prior findings elhage2021mathematical; dong2021attention that biases encode “default” priors and are crucial for rank restoration, allowing us to shift reasoning dispositions without corrupting the semantic correlations stored in weights (see more discussion in Section 4.2).
+Our decision to target bias-like offsets is motivated by prior findings Elhage et al.
+(2021); Dong et al.
+(2021) that biases encode “default” priors and are crucial for rank restoration, allowing us to shift reasoning dispositions without corrupting the semantic correlations stored in weights (see more discussion in Section 4.2).
 Concretely, for a hidden state $\mathbf{x}$, we implement:
-$\mathbf{y}=\text{Norm}(\mathbf{x})\cdot\mathbf{w}+(\mathbf{b}+\mathbf{z}),\quad\mathbf{z}\sim\mathcal{N}(\mu,\sigma^{2})$ (4)
-Here, $\mu$ and $\sigma$ are hyperparameters controlling the center and spread of the population diversity, and $\mathbf{b}$ can be 0 in RMSNorm.
-For normalization variants that include a learned scale (and optionally a bias), these parameters remain part of the deterministic backbone, while $\mathbf{z}$ provides a stochastic post-normalization offset.
-Sampling $\mathbf{z}$ therefore yields multiple model instances that share the same pre-trained weights but differ in their normalization regime.
-4.2 Biases as Dispositional Priors.
-B-Trans introduces stochasticity through additive offsets rather than through full weight perturbations.
+$\mathbf{y}=\mathrm{Norm}(\mathbf{x};\mathbf{w},\mathbf{b})+\mathbf{z},\quad\mathbf{z}\sim\mathcal{N}(\boldsymbol{\mu},\,\sigma^{2})$ (4)
+Here, $\mu$ and $\sigma$ are hyperparameters controlling the center and spread of the population diversity, and $\mathbf{b}$ can be 0 in RMSNorm (which makes $z$ a post-norm additive offset).
+For normalization variants that include a learned scale (and optionally a bias), these parameters remain part of the deterministic backbone, while $\mathbf{z}$ provides a stochastic post-normalization offset for each sampled model instance.
+4.2 Biases as Dispositional Priors
+B-Trans introduces stochasticity through additive offsets rather than through full weight modification.
 This choice is grounded in the functional role of additive shifts in Transformer mechanics.
-First, additive components act as input-independent defaults: while weight matrices encode correlations between features, additive shifts set baseline writes to the residual stream and influence activation thresholds before input-specific computation elhage2021mathematical.
+First, additive components act as input-independent defaults: while weight matrices encode correlations between features, additive shifts set baseline writes to the residual stream and influence activation thresholds before input-specific computation Elhage et al.
+(2021).
+Perturbing these terms, therefore, alters a model’s inductive dispositions without corrupting the semantic relationships stored in its weights.
 Second, additive vectors are a practical handle for semantic steering.
-Work on activation engineering shows that adding vectors in the residual stream can steer high-level behaviors without broadly degrading capabilities turner2023steering; zou2023representation.
-Third, bias-like shifts can be important for defining operating points.
-Recent evidence on massive activations suggests that when explicit biases are absent, models may simulate similar effects through outlier activations sun2024massive.
-By sampling $\mathbf{z}$, B-Trans creates a population of dispositions that explore different reasoning trajectories while sharing the same underlying knowledge in the backbone weights.
+Works on activation engineering show that adding vectors in the residual stream can steer high-level behaviors without broadly degrading capabilities Turner et al.
+(2023); Zou et al.
+(2023).
+Third, bias-like shifts can be critical for model behavior.
+Sun et al.
+(2024) suggests that when explicit biases are absent, models may simulate similar effects through outlier activations.
+For these reasons, B-Trans targets additive offsets as a minimal and controlled locus for inducing population diversity.
+While extending stochasticity to additional parameters—such as normalization weights $w$ (Equation 4)—may further enrich diversity, we do not explore such extensions in this work.
+Perturbing full weight matrices of LLMs, in contrast, risks disrupting learned representations and introducing excessive noise, potentially degrading model behavior.
+We therefore restrict our study to additive offsets and leave broader normalization- or weight-level stochasticity to future work.
 4.3 Temporal Consistency via Noise Caching
-A requirement for valid variational approximation in sequential models is maintaining a consistent stochastic mask across time steps, as established in prior work on Recurrent Neural Networks gal2016dropout.
-Naively sampling $\mathbf{z}$ at every token step violates this principle, resulting in a temporally incoherent model that effectively shifts its functional hypothesis mid-sentence.
-To adhere to this theoretical constraint within our framework, we implement a sequence-level noise caching mechanism (see Figure 1).
-As detailed in the BayesianBiasWrapper, the process ensures that the stochastic perturbation defines a stable reasoning persona for the duration of the generation: (1) Initialization: Upon receiving the prompt, we sample the latent variable $\mathbf{z}$ once and cache it, effectively drawing a single sample $\theta^{(k)}$ from the approximate posterior.
+As mentioned in Section 3.2, for each sequence generation, we sample one LLM (analogous to a distinct individual from the population) to solve the problem.
+Thus, we implement a sequence-level noise caching mechanism (see Figure 1).
+As detailed in the BayesianBiasWrapper, the process ensures that the stochastic perturbation defines a stable reasoning flow for the entire response: (1) Initialization: Upon receiving the prompt, we sample the latent variable $\mathbf{z}$ once and cache it, effectively drawing a single sample $\theta^{(k)}$ from the approximate posterior.
 (2) Generation: For all subsequent autoregressive steps within the same sequence, the model applies this frozen $\mathbf{z}$ to the normalization layers.
@@
 Crucially, unlike full weight-masking methods (e.g., MC Dropout), our implementation is computationally negligible, requiring only an element-wise addition of a cached tensor without the overhead of generating large binary masks.
-Specifically, implementing theoretically correct (frozen) MC Dropout on a 7B model would require caching around 14GB of masks per user instance, assuming pre-expand the binary masks to FP16 to align with weight precision to avoid prohibitive on-the-fly decompression overheads.
+Specifically, implementing theoretically correct (frozen) MC Dropout on a 7B model would require caching around 14GB of masks per user instance, assuming pre-expanding the binary masks to FP16 to align with weight precision to avoid prohibitive on-the-fly decompression overheads.
 In contrast, B-Trans only caches the bias noise vectors (<1MB), reducing the memory footprint by orders of magnitude while maintaining temporal consistency.
 5 Experiments
-Our evaluation probes B-Trans from three complementary angles.
-First, we test whether sampling model instances produces functional diversity in zero-shot generation rather than surface-level randomness.
-Second, we study exploration in Reinforcement Learning with Verifiable Rewards (RLVR), asking whether parameter-space sampling improves search under sparse rewards.
-Third, we examine population aggregation in a label-free Test-Time RL (TTRL) setting, where the learning signal is induced from model outputs rather than ground-truth supervision.
-Across all settings, our goal is to provide a minimal proof-of-concept for population-based exploration under controlled, standard protocols, rather than an exhaustive benchmark.
-Accordingly, we keep training and decoding setups aligned with established baselines and focus on whether B-Trans yields consistent improvements in exploration and semantic diversity.
-It is important to note that while B-Trans modifies the architecture’s normalization layers, the modified backbone remains differentiable and trainable via standard backpropagation pipelines.
-From the efficiency and computing concern, we follow Low-Rank Adaptation (LoRA) hu2022lora training for the RLVR and TTRL tasks.
-In all training settings, we inject stochasticity only during rollout for exploration, while keeping the update phase deterministic by using the mean shift only, matching standard optimization stability requirements.
+To validate that B-Trans can produce diverse but meaningful minds from one model, we conduct experiments under two settings: (1) B-Trans for zero-shot generation, focusing on improving semantic diversity and also pass@k for accuracy, and (2) B-Trans for output sampling, focusing on providing a useful learning signal for Reinforcement Learning with Verifiable Rewards (RLVR).
+Across both settings, our goal is to provide a minimal proof-of-concept for population-based exploration under controlled, standard protocols, rather than an exhaustive benchmark.
+5.1 Zero-Shot Diversity and Creativity
+The premise of B-Trans is that each sampled response corresponds to a different model instance.
+We evaluate whether B-Trans yields meaningful heterogeneity compared to the deterministic baseline with high-temperature decoding.
 Figure 2: Pass@$K$ performance on MMLU-Pro across Llama-3.1 and Qwen3 families.
@@
 Figure 4: PCA visualization of response embeddings from Qwen3 and Llama-3.1 models.
-The query is “Output a hard question to humanity (super concise and short), independent of theme.” Standard sampling confines models to disjoint semantic regions (their respective “territories”).
-B-Trans significantly expands this scope, allowing the model to escape its default mode and even explore semantic regions typically associated with the opposing model (highlighted in red box).
-5.1 Zero-Shot Diversity and Creativity
-The premise of B-Trans is that different instance samples correspond to distinct functional behaviors.
-We evaluate whether B-Trans yields meaningful heterogeneity compared to high-temperature decoding.
+The query is “Output a hard question to humanity (super concise and short), independent of theme.” Standard sampling confines models to small semantic regions.
+B-Trans significantly expands the territories, allowing the model to escape its default mode and even explore semantic regions typically associated with the other models (highlighted in red box).
 Setting.
 We study intrinsic population behavior on two benchmarks with different measurement goals.
-On MMLU-Pro wang2024mmlu (reasoning), we use Pass@K to test whether additional sampled individuals provide functional and valid reasoning paths: for each question we sample $K$ independent sequence-level personas and count the answer as correct if any sample is correct.
-Thus, improved scaling with larger $K$ indicates that the induced diversity is not mere noise but yields additional correct solutions.
-On INFINITY-CHAT jiang2025artificial (open-ended creativity), where there is no ground-truth correctness, we measure semantic diversity using average pairwise embedding cosine distance reimers2019sentence; song2020mpnet and visualize the embedding geometry with PCA following Artificial Hivemind (jiang2025artificial).
-The goal is to quantify and visualize semantic heterogeneity under the same generation setup as generating answer for MMLU-Pro.
-For both tasks, we contrast B-Trans against two distinct base models, Qwen3 yang2025qwen3 and Llama-3.1 grattafiori2024llama, using high-temperature sampling ($T=1$).
-This setup allows us to observe whether the models can escape their default semantic modes through B-Trans versus Temperature stochasticity.
+On MMLU-Pro Wang et al.
+(2024) (a challenging scientific reasoning dataset that contains 12K complex questions across various disciplines), we use Pass@K to test whether additional sampled individuals provide valid reasoning paths: for each question, we sample $K$ independent LLM instances and count the answer as correct if any sample is correct.
+Thus, improved scaling with larger $K$ indicates that the induced diversity is not mere repetition but yields a correct solution.
+On INFINITY-CHAT Jiang et al.
+(2025) (100 open-ended questions that admit a wide range of plausible answers with no ground-truth correctness), we measure semantic diversity using average pairwise embedding cosine distance Reimers and Gurevych (2019); Song et al.
+(2020) and visualize the embedding geometry with PCA following Artificial Hivemind analysis (Jiang et al., 2025).
+The goal is to quantify and visualize semantic heterogeneity.
+For both tasks, we contrast B-Trans against two distinct base models, Qwen3 1.7B, 8B Yang et al.
+(2025), and Llama-3.1 Grattafiori et al.
+(2024).
+All models share the same decoding strategies.
 Results.
 On MMLU-Pro (Figure 2), while B-Trans matches the baseline at $K=1$, it demonstrates significantly stronger scaling behavior as the sample budget increases.
-This trend confirms that the diversity introduced by B-Trans is semantically meaningful: rather than generating random noise (which would yield diminishing returns), the implicit population successfully explores valid reasoning trajectories that are inaccessible to the deterministic model.
-Complementing this, the embedding distance analysis on INFINITY-CHAT (Figure 3) further quantifies this behavior, confirming superior output heterogeneity.
-Qualitatively, Figure 4 visualizes the semantic landscape.
-Under standard sampling, Qwen3 and Llama-3.1 occupy disjoint, narrow regions in the semantic space, reflecting their distinct pre-trained biases.
-While high temperature slightly expands these clusters, it fails to alter their fundamental topology.
+This trend confirms that the diversity introduced by B-Trans is semantically meaningful: rather than generating redundant repetition (which would yield diminishing returns), the B-Trans population successfully explores valid reasoning trajectories that are not found by the deterministic baselines.
+Complementing this, the embedding distance analysis on INFINITY-CHAT (Figure 3) further quantifies this behavior, confirming better output heterogeneity.
+Qualitative Analysis
+Figure 4 visualizes the semantic landscape, where spatial proximity indicates the semantic similarity of response embeddings (dots).
+Note that the horizontal distance reflects the most fundamental conceptual differences between each response.
+Under standard sampling (temp-0.5), Qwen3 and Llama-3.1 occupy disjoint, narrow regions in the semantic space, reflecting their distinct pre-trained biases.
+While high temperature (temp-1) slightly expands these clusters, it fails to alter their fundamental topology.
 In contrast, B-Trans generates a significantly broader distribution, forming distinct clusters that explore diverse reasoning paths.
-Analysis.
-The visualization in Figure 4 reveals a critical insight into the nature of the generated diversity.
-Standard temperature sampling is confined to local perturbations; it cannot push a model out of its inherent “personality,” leaving Qwen3 and Llama-3.1 isolated in their respective semantic territories.
-Conversely, B-Trans enables the models to transcend these native constraints.
-The visualization shows that B-Trans samples not only cover a wider area but also begin to bridge the gap between the two models, with samples from one model occasionally reaching the semantic territory typically occupied by the other.
-This suggests that perturbing normalization biases can shift the model’s generation distribution in embedding space beyond local temperature-based perturbations.
-Meanwhile, the overlap indicates that B-Trans can occasionally reach semantic regions that are atypical for the original base model under the same decoding budget, thus overlapping with the reasoning styles of different model families, rather than merely adding noise to a single persona.
+Notably, B-Trans can occasionally reach semantic regions that are atypical for the original base model.
+As examples highlighted in the red box, Qwen3-8B-btrans generates one response that is very close to Llama-3.1’s style.
+This validates our goal of mining different minds from one model.
 5.2 Exploration in Reinforcement Learning with Verifiable Rewards (RLVR)
+In Reinforcement Learning, the critical bottleneck is often the exploration-exploitation trade-off.
+Standard LLM policies rely on action-space noise (sampling tokens) to explore.
+In this section, we show that parameter-space noise (sampling models) provides a more consistent and deeper form of exploration under the setting of RLVR.
+Setting.
+We follow the default configuration of the SimpleRL-Zoo framework as-is, which is a popular RLVR recipe to improve models’ reasoning abilities using Group Relative Policy Optimization (GRPO) Guo et al.
+(2025).
+Our studies include different sizes of Qwen3 models (1.7B and 8B), math-specialized Qwen2.5-Math-7B, and Llama3.1-8B.
+For each model, we compare two approaches to sample a group of responses (16 outputs) for each input for GRPO reward computing : (1) using a deterministic base model, and (2) using B-Trans.
+Note that we keep the policy model the same for both settings and focus solely on the difference in sampling strategies.
+To save computing, we only update Low-Rank Adaptation (LoRA) Hu et al.
+(2022) adapters for the policy model instead of full-parameter fine-tuning (less than 1% trainable parameters), and we train on 400 examples sampled from the same distribution of the SimpleRL-Zoo dataset via difficulty-stratified sampling.
+We test the finetuned model on three standard mathematical reasoning datasets, i.e., GSM8K Lewkowycz et al.
+(2022), MATH-500 Lightman et al.
+(2023), and Minerva Math Lewkowycz et al.
+(2022), where the ground truth is verifiable.
+Results.
+As shown in Table 1, B-Trans achieves consistent gains across all four benchmarks.
+GRPO relies on informative relative comparisons within the sampled group.
+When the rollout group is homogeneous, relative advantages become less informative, and updates can stagnate.
+By producing group members from distinct parameter-space instances, B-Trans increases the chance that at least one candidate reaches a higher-reward region, strengthening the learning signal.
 Model GSM8K MATH-500 Minerva Math Average
@@
       $\hookrightarrow$ B-Trans 78.6 23.7 9.9 37.4
-Table 1: Performance on RLVR tasks using GRPO. B-Trans consistently outperforms the baseline, particularly on smaller models and harder datasets.
-In Reinforcement Learning, the critical bottleneck is often the exploration-exploitation trade-off.
-Standard LLM policies rely on action-space noise (sampling tokens) to explore.
-We hypothesize that parameter-space noise (sampling models) provides a more consistent and deeper form of exploration.
-By solving mathematical problems where the solution space is sparse, we test the competency of the implicit B-Trans population.
-If B-Trans achieves higher rewards under Group Relative Policy Optimization (GRPO) guo2025deepseek, it indicates that the diverse rollout candidates are not just different, but are hitting the correct solution regions more frequently than the baseline.
-Setting.
-To demonstrate the impact of B-Trans on exploration and exploitation trade-offs, we utilize the standard RLVR setting, following the SimpleRL-Zoo framework  [^2] We follow the default configuration as-is; the only change is apply BayesianBiasWrapper during the rollout, as we do not introduce any additional modifications beyond B-Trans..
-To save computing, we use LoRA instead of full-parameter fine-tuning (less than 1% trainable parameters), and we train on 400 examples sampled from the same distribution of SimpleRL-Zoo dataset via difficulty-stratified sampling.
-In GRPO, a group of outputs is sampled for each input, and the policy is updated based on their relative rewards.
-We test on four standard mathematical reasoning datasets, i.e., GSM8K lewkowycz2022solving, MATH-500 lightman2023lets, and Minerva Math lewkowycz2022solving, where the ground truth is verifiable.
-Results.
-We compare GRPO training with standard decoding against GRPO training where B-Trans instance sampling is enabled during rollout, under the same protocol.
-As shown in Table 1, B-Trans achieves consistent gains across all four benchmarks.
-GRPO relies on informative relative comparisons within each sampled group.
-When the rollout group is homogeneous, relative advantages become less informative and updates can stagnate.
-By producing group members from distinct parameter-space instances, B-Trans increases the chance that at least one candidate reaches a higher-reward region, strengthening the learning signal available to GRPO under the same group size.
-5.3 Test-Time RL without Labels
-This experiment represents the most challenging scenario: learning without ground-truth labels.
-We employ Test-Time RL (TTRL), a setting where the model must self-improve by generating candidate solutions and filtering them via consistency checks.
-Specifically, we follow the established TTRL protocols zuo2025ttrl; yu2025restrain to utilize majority voting as the reward proxy.
-It is worth noting that the majority voting inherently rewards consensus, seemingly contrasting with B-Trans’s philosophy of encouraging diverse exploration (“The truth is often with the few”).
-However, our contribution focuses on the quality of the exploration rather than the selection mechanism.
-We argue that a superior learning signal should emerge from diverse independent reasoning paths rather than simple outcome consensus from mode collapse.
-Under the GRPO framework, standard models often generate homogeneous samples, providing redundant gradient information even when the reward is positive.
-Conversely, B-Trans injects parameter-space diversity, ensuring that the high-reward trajectories (those aligning with the majority) exhibit varied reasoning paths.
-This semantic diversity maximizes the information gain per update step, allowing the model to explore diverse reasoning patterns rather than overfitting to a single, narrow solution path.
-Setting.
-We investigate the efficacy of B-Trans in scenarios characterized by sparse signals and challenging search spaces using the AIME24 dataset.
-Unlike standard RLHF, TTRL adapts the model using only unlabeled test prompts.
-To synthesize a reward signal without ground truth, we utilize an outcome-based Majority Voting mechanism.
-For each input, we sample $N$ candidate responses and extract the final numerical answer.
-A consensus label is determined solely by the most frequent final answer.
-Trajectories leading to this consensus answer are assigned a positive reward ($r=1$), while others receive zero.
-Crucially, this reward ignores the semantic diversity or logical validity of the intermediate reasoning steps (Rationale), making it a coarse signal that relies heavily on the diversity of the generated candidates to be effective.
-All parameter updates are applied via LoRA for computational efficiency.
-Results.
-Figure 5 illustrates the reward curves evaluated every 5 steps.
-We compare the standard Qwen3-1.7B model using high-temperature sampling (temp=1.0) against B-Trans.
-The B-Trans exhibits high variance and oscillation but achieves significantly higher peak accuracy compared to the baseline.
-Figure 5: Accuracy curve during training. Reported at every five training steps.
-Analysis.
-Two key insights emerge from this experiment.
-First, the hyperparameter $\sigma$ serves as an effective control for exploration intensity.
-Higher values allow the model to escape local optima through aggressive exploration, akin to brainstorming wild ideas, which is necessary for the challenging AIME24 problems.
-Second, the success of B-Trans in the TTRL setting suggests an emergent self-correction capability.
-This parallels human cognitive processes, where internal deliberation and re-evaluation can refine understanding even without external feedback.
-Even with a coarse consensus-based reward, B-Trans effectively filters out homogeneous cognition common to single-model instances by providing a broader set of rationales for the vote.
-6 Ablation Study and Discussion
-In this section, we dissect the architectural choices of B-Trans to validate our core design principles.
-We systematically investigate the necessity of temporal consistency in noise injection, justify the decision to perturb biases over weights, and verify the robustness of our approach across different post-training frameworks.
-6.1 Temporal Consistency: Sequence-wise vs. Token-wise Noise
-To validate the necessity of hypothesis sampling (Section 3.2), we compare B-Trans against a Token-wise Noise baseline on MATH-500 using Qwen3-1.8B.
+Table 1: Accuracy (%) on RLVR tasks using GRPO. B-Trans consistently outperforms the baseline.
+6 Ablation Study
+To validate the necessity of hypothesis sampling (Section 3.2), we compare B-Trans against a Token-wise Noise baseline on MATH-500 using Qwen3-1.7B.
 While B-Trans samples a single latent $z$ per sequence, the baseline resamples $z_{t}\sim\mathcal{N}(0,\sigma^{2}I)$ at every forward pass.
-Crucially, we fix an identical noise magnitude ($\sigma=0.02$) for both settings to isolate the impact of temporal inconsistency.
+We fix an identical noise magnitude ($\sigma=0.02$) for both settings to isolate the impact of temporal inconsistency.
 We use the Step-wise Consistency Score (SCS) to quantify the logical stability of the reasoning chain.
@@
 $\text{SCS}(Y)=\frac{1}{m-1}\sum_{i=1}^{m-1}\cos(\text{Enc}(s_{i}),\text{Enc}(s_{i+1}))$ (5)
-Results & Analysis.
+Results and Analysis
 The Token-wise baseline suffers catastrophic degradation.
@@
 46.4%).
-This empirical failure corroborates the theoretical insight established by gal2016dropout at RNNs/LSTMs age: valid variational inference requires the stochastic mask to remain invariant across time steps to represent a single coherent hypothesis.
-B-Trans satisfies this “Temporal Consistency” requisite through lightweight bias perturbation, effectively capturing the benefits of weight uncertainty without the memory overhead of full ensemble masks.
-6.2 Post-training Infrastructure Agnosticism
-To validate that B-Trans serves as a "plug-and-play" module—agnostic not just to the model architecture but also to the training infrastructure, we repeat our TRL-based vonwerra2022trl RLVR experiments (Section 5.2) using another widely adopted RL engine, namely VeRL sheng2024hybridflow (Volcano Engine Reinforcement Learning).
-We observed no statistically significant difference in the final convergence performance or learning dynamics between the two implementations.
-While the VeRL implementation exhibited faster wall-clock training speeds in our environment, we attribute this to differences in rollout acceleration libraries rather than the B-Trans mechanism itself.
-We expect B-Trans to be agnostic to the underlying base LLMs and training infrastructure, supporting its potential as a surgical plug-and-play approach for diverse use cases.
-7 Future Work
+This empirical failure corroborates the theoretical insight established by Gal and Ghahramani (2016) at RNNs/LSTMs age: valid variational inference requires the stochastic mask to remain invariant across time steps to represent a single coherent hypothesis.
+B-Trans satisfies this Temporal Consistency requisite through lightweight bias perturbation, effectively capturing the benefits of weight uncertainty without the memory overhead of full ensemble masks.
+7 Future Work and Discussion
 Our current implementation models the prior distribution as a simple isotropic Gaussian with a scalar hyperparameter $\sigma$.
 While this serves as an effective proof-of-concept, the assumption of uniform uncertainty across all layers and dimensions is a strong simplification.
-A natural progression for this line of research is to develop a learnable or hierarchical prior.
-In this initial exploration, we adhered to a fixed prior to mitigate the risks associated with post-training optimization.
+Furthermore, as mentioned in Section 4.2, we could also introduce a distribution over parameters other than Bias-like offset, and the weights in normalization could be a natural target.
+Another progression for this line of research is to develop a learnable or hierarchical prior.
+In this initial exploration, we adhered to a fixed $\sigma$ to mitigate the risks associated with post-training optimization.
 Learning variance parameters on limited downstream datasets can lead to overfitting, where the induced population loses its general competence.
+Ideally, the population model can be learned from a large corpus reflecting real population behavior, similar to using real human interaction data for Bayesian recommender system training Zhang and Koren (2007).
 We anticipate that future efforts, particularly those with sufficient computational resources to train on large-scale general corpora, could successfully learn these uncertainty parameters.
-Such a data-driven prior would likely act as a more robust “population,” capturing layer-specific sensitivities that a scalar $\sigma$ cannot.
+Such a data-driven prior would likely act as a more robust population, capturing layer-specific sensitivities that a scalar $\sigma$ cannot.
 Furthermore, the scalar $\sigma$ presents an opportunity for adaptive inference.
@@
 8 Conclusion
-We revisit the prevailing view of Large Language Models as static point estimates and argue that lightweight parameter-space uncertainty can be practically useful for exploration and diversity.
-We introduce B-Trans, a population-based framework that samples diverse yet coherent model instances from a single Transformer backbone using a variationally motivated Gaussian posterior proxy.
-By applying temporally consistent perturbations to normalization biases, B-Trans simulates the wisdom of crowds at negligible cost, yielding significant gains in zero-shot diversity and sparse-reward reinforcement learning.
-Our results suggest that representing uncertainty as a sampleable local proxy, even in a simple form, can enable population-driven reasoning within a single pre-trained model.
+We revisit the prevailing view of Large Language Models as static point estimates and argue that lightweight parameter-space modification is a practical catalyst for exploration and diversity.
+Benefited from the wisdom of crowds, our experiments demonstrate that B-Trans, even with a simple instantiation, provides meaningful and diverse responses at negligible cost.
+In an era where foundation models have achieved strong deterministic prediction capabilities (exploitation), we envision that B-Trans contributes to the next generation of models by breaking through structural homogenization and facilitating efficient learning via deeper exploration.
 Limitations
 While B-Trans demonstrates a promising direction for inducing implicit ensembles, our current study is intentionally scoped as a minimal, representative proof-of-concept centered on architectural simplicity and controlled comparisons.
-We do not perform extensive hyperparameter sweeps, broad architecture coverage, or large scale benchmarking beyond validating the key design principles.
-Our intervention is restricted to the bias terms of normalization layers, which is a simplified proxy rather than a full Bayesian neural network.
+We do not perform extensive hyperparameter sweeps, broad architecture coverage, or large-scale benchmarking beyond validating the key design principles.
+Our intervention is restricted to the biased terms of normalization layers, which is a simplified proxy rather than a full Bayesian neural network.
 While this surgical approach is computationally negligible, it is a simplified approximation of a full Bayesian Neural Network.
 The weights in the attention and feed-forward layers remain deterministic, potentially limiting the expressivity of the induced posterior.
+Ethics Considerations and Broader Impact
+This work introduces a parameter-space sampling method to enhance LLM diversity during inference.
+Regarding potential risks, we acknowledge that techniques improving generative diversity could theoretically be misused to generate varied disinformation or spam.
+However, our primary objective is to enable deeper exploration in reasoning tasks.
+Environmentally, B-Trans offers a positive impact by approximating population-level intelligence from a single model, thereby avoiding the significant carbon footprint associated with training and maintaining full model ensembles.